# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v6

@@ 4.2 Risk-coping Mechanisms
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity. [^60] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I prefer to use the conservative setting to deal with potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. As for $r_{i,t}$, I consider several variables to overview various risk-coping strategies. These include the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as the difference between income and expenses. The estimates of $\kappa$ are therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also consider the earnings of all the family members except for the head and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blundell et al. 2016). [^61] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focuses on the households living in Tsukishima throughout the sample period, I can discard addressing the ex-ante risk-coping strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables are expected to be negative, reflecting additional earnings when the head’s income is reduced.
@@ Results
Net savings -0.043 [0.018]** 289 -0.042 [0.018]* 162
Net insurance 0.001 [0.013] 289 0.015 [0.024] 162
Net borrowing -0.021 [0.027] 289 0.036 [0.036] 162
Net credit purchase -0.128 [0.039]*** 289 -0.017 [0.059] 162
Net gifts -0.052 [0.019]** 289 -0.015 [0.014] 162
@@
Other member’s earnings 0.001 [0.010] 289 0.038 [0.024] 162
Sales of assets -0.000 [0.000] 289 0.000 [0.000] 162
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns are close to zero and statistically insignificant. Although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^64] The OISR, Archives of the Tsukishima Survey (THBS #1, 13th January 1919). In addition, I find that the sales of miscellaneous assets such as newspapers, bottles, and clocks did not work as the risk-coping.
@@ 4.3 Event-Study Analysis: Response to the Advers Health Shocks
Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture between January 1918 and December 1920. The death statistics used in this figure were collected based on the comprehensive registration system. Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
Figure 5: Number of Influenza Deaths in Tokyo between 1918 and 1920
@@
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings shall mainly be influenced by the number of days suffering illness in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). Thus, this may reflect the reductions in their salary due to the absence. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^66] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income. Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
Table 6: Results of Testing the Impacts of Adverse Health Shockson the Head’s Earning
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including eather of the two variables does not change the finding on the neglisible impacts of the health shocks on the other family members.
@@
where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system.
@@
Table 7: Results of Estimating Income Elasticities:Smoothing Consumption against the Adverse Health Shocks
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects. The estimated coefficients on the health shock variable are listed in the sub-column named “Coef.”. Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficients on the disposable income are listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 in the spirit of Cochrane (1991). Column 1 in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting the evidence that adverse health shocks on the heads reduce consumption levels. The estimate for the food subcategory is relatively small. The estimates for the clothing and gifts categories show a relatively large magnitude. These are in line with my baseline results in Table 4. The estimate for the education category is now statistically significantly negative. This may reflect the reduction of allowances for children in response to income losses due to health shocks.
@@
Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3. If the negative impacts of the health shocks on consumption were derived from the income losses, the result from the structural estimation approach is similar to that in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. The results are materially similar to the reduced-form results in Column (1). The estimates are systematically larger than those reported in Table 4. While this result shall reflect the bias reduction thanks to the instrumental variable estimation, it is also important to remember that this system is designed to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events. [^67] In other words, the furniture may have been purchased only when the households had extra income. This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^68] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applicable to the transportation category because the family travels did not happen frequently. These cases may be another explanation for the difference in the results from my baseline results in Table 4.
@@
While the estimates for clothes and gifts categories now exceed one, the null of equality to the population mean of one is not rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on the rare health shock events, the findings are materially similar to my baseline results: both categories are sensitive to the shocks as those are dispensable categories.
@@
Table 8: Results of Testing the Risk-Coping Mechanisms:Adverse Health Shock as the Idiosyncratic Shock
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the subcolumn named “Coef.”. Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earning are listed in the subcolumn named “Coef.”. Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.

### v7

@@ 4.2 Risk-coping Mechanisms
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity. [^64] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I prefer to use the conservative setting to deal with potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. As for $r_{i,t}$, I consider several variables to overview various risk-coping strategies. These include the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as the difference between income and expenses. The estimates of $\kappa$ are therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also consider the earnings of all the family members except for the head and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blundell et al. 2016). [^65] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focuses on the households living in Tsukishima throughout the sample period, I can discard addressing the ex-ante risk-coping strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income is reduced.
@@ Results
Net savings -0.043 [0.018]** 289 -0.035 [0.028] 124
Net insurance 0.001 [0.013] 289 0.002 [0.033] 124
Net borrowing -0.021 [0.027] 289 -0.027 [0.017] 124
Net credit purchase -0.128 [0.039]*** 289 -0.011 [0.042] 124
Net gifts -0.052 [0.019]** 289 -0.006 [0.029] 124
@@
Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
Sales of assets -0.000 [0.000] 289 0.000 [0.001] 124
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes:
This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns are close to zero and statistically insignificant. Miyoshi (1989, pp. 383–386) documented that the wives regularly engaged in side jobs rather than worked temporarily to compensate for the income loss. In fact, although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^68] The OISR, Archives of the Tsukishima Survey (THBS#1, 13th January 1919). In the final row in Panel B, I find that selling miscellaneous assets such as newspapers, bottles, and clocks did not work as a risk-coping strategy.
@@ 4.3 Event-Study Analysis: Response to the Advers Health Shocks
Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture between January 1918 and December 1920. The death statistics used in this figure were collected based on the comprehensive registration system.
Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
@@
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings shall mainly be influenced by the number of days suffering illness in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). Thus, this reflects the reductions in their salary due to the absence: Miyoshi (1989, pp. 383; 388) highlighted that income was accidentally reduced due to absence when the head was infected with a disease. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^70] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income. Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
Table 6: Results of Testing the Impacts of Adverse Health Shocks
on the Head’s Earning
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including eather of the two variables does not change the finding on the neglisible impacts of the health shocks on the other family members.
@@
where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. [^71] As noted in Section 2.3, the factories did not offer compensation for their sickness leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916). Even in the largest factory in Tsukishima, the Ishikawajim Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930). In fact, Miyoshi (1989) documented the issue that the head’s disease directly causes income loss due to the absence. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system.
@@
Table 7: Results of Estimating Income Elasticities:
Smoothing Consumption against the Adverse Health Shocks
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes:
@@
Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects. The estimated coefficients on the health shock variable are listed in the sub-column named “Coef.”.
Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficients on the disposable income are listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
@@
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 in the spirit of Cochrane (1991). Column (1) in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels. Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3. If the negative impacts of the health shocks on consumption were derived from the income losses, the result from the structural estimation approach is similar to that in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. As shown, the results are materially similar to the reduced-form results in Column (1). Importantly, the estimates are systematically larger than those reported in Table 4. The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$. While this result partly reflects the bias reduction under the instrumental variable estimation, it is also important to remember that this system is designed to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events. [^72] In other words, the furniture may have been purchased only when the households had extra income. This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^73] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applicable to the transportation category because family travel is not a frequent event and thus, unlikely to coincide with the timing of illness. In contrast, the entertainment expenses are still likely to be cut in case of illness, which may reflect the curtailment of tobacco, theater-going, and the allowance for heads. This may be another explanation for the difference in the results from my baseline results in Table 4.
@@
While the estimates for clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on the rare health shock events, the findings are materially similar to my baseline results: both categories are sensitive to the shocks as those are dispensable categories.
@@
Table 8: Results of Testing the Risk-Coping Mechanisms:
Adverse Health Shock as the Idiosyncratic Shock
@@
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes:
Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the subcolumn named “Coef.”.
Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earning are listed in the subcolumn named “Coef.”.
Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.